# Graph Conditional Flow Matching for Relational Data Generation

### v1

@@ 1 Introduction
Tabular data synthesis [48, 49, 27, 50, 22] has been subject to research for several years. Early methods were based on Bayesian networks [51], factor graphs [33] and autoregressive models [34]. Most recent methods leverage the breakthroughs of deep-learning based generative models, from early latent variables models as VAEs [26] and GANs [13], to transformer-based autoregressive models [45] and diffusion models [19].
@@
Despite the advancements in single table synthesis, generating multiple tables characterized by foreign-key constraints is a considerably more difficult task. Relational datasets can be represented as large graphs, where nodes correspond to records and edges denote foreign-key relationships. This introduces a dual challenge: (1) modeling potentially complex graph structures such as tables with multiple parents, or multiple types of relationships between two tables and (2) modeling statistical dependencies between records linked directly or indirectly through foreign keys.
@@
Relational data generation [38, 15, 43, 36] is a less mature field, with few existing methods capable of properly handling complex database structures. In Xu et al. [46], they separately model the graph structure and then propose a method for generating the content of the relational database table-by-table. The generation of each record is conditioned on graph-derived node statistics and aggregated information from connected records. In concurrent work, Hudovernik [20] follow a similar approach, generating the content of the tables using a latent diffusion model conditioned on node embeddings produced by a separate model.
@@
In this work, we propose a method for generating the content of a relational dataset given the graph describing its structure. Inspired by recent advancements in image generation, we employ flow-matching to train a flow-based generative model of the content of the entire relational dataset. In order to enable information propagation across connected records, the architecture of the learned denoiser includes a graph neural network (GNN) [41, 2]. This approach aims at maximizing expressiveness in modeling correlation between different records of the database, as information can be passed arbitrarily within a connected component through a GNN. Moreover, our framework is flexible as the conditioning graph can be complex, and scalable as we can generate large datasets.
@@
Using SyntheRela [21], a recently developed benchmarking library, we prove the effectiveness of our method on several datasets, comparing it with several open-source approaches. Experimental results show our method achieves state-of-the-art performance in terms of fidelity of the generated data.
Our implementation is publicly available. [^1] Code: https://github.com/DavideScassola/graph-conditional-flow-matching
Figure 1: Overview of the architecture of the denoiser for relational data. A relational dataset composed of multiple tables can be seen as a graph, where records are the nodes and foreign keys are the edges. The denoiser takes as input a relational dataset where noise was added to each record with noise level $t$. Firstly, a graph neural network (GNN) processes the entire graph and computes node embeddings $\varepsilon_{i}$ encoding context information for each record. Each record and its corresponding embedding are then processed independently by table-specific multi-layer perceptrons (MLPs), which predict the original clean records ($t=1$).
@@ 2.1 Continuous normalizing flows
Flow matching [29] is an emerging framework for training continuous normalizing flows (CNFs), a deep generative model that learns to transform random noise into target distributions through ordinary differential equations.
@@ 2.2 The flow matching objective
Since direct access to the vector field $v_{t}$ of a CNF generating the data is unavailable, we cannot directly match a parameterized velocity field $v_{t}^{\theta}$ against the ground truth $v_{t}$. The main idea of flow matching is instead to define the underlying probability path as a mixture of conditional "per-example" probability paths, that can be defined in a tractable way.
Let us denote by $p_{t}(\boldsymbol{x}|\boldsymbol{x}_{1})$ a conditional probability path such that $p_{0}(\boldsymbol{x}|\boldsymbol{x}_{1})=p_{0}(\boldsymbol{x})$ and $p_{1}(\boldsymbol{x}|\boldsymbol{x}_{1})$ is a distribution concentrated around $\boldsymbol{x}=\boldsymbol{x}_{1}$, as $p_{1}(\boldsymbol{x}|\boldsymbol{x}_{1})=\mathcal{N}(\boldsymbol{x}|\boldsymbol{x}_{1},\sigma^{2}I)$ with $\sigma$ small. We define the conditional velocity $u_{t}(\boldsymbol{x}\mid\boldsymbol{x}_{1})$ as the velocity generating the conditional probability path $p_{1}(\boldsymbol{x}|\boldsymbol{x}_{1})$. It is then possible to prove that the marginal velocity, defined as
[…]
@@
$\mathcal{L}_{\mathrm{FM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}[0,1],\boldsymbol{x}_{1}\sim q(\boldsymbol{x}_{1}),\boldsymbol{x}\sim p_{t}(\boldsymbol{x}\mid\boldsymbol{x}_{1})}\|v_{t}^{\theta}(\boldsymbol{x})-u_{t}(\boldsymbol{x}\mid\boldsymbol{x}_{1})\|^{2}$
known as the conditional flow matching objective, where a neural network learns a match the marginal velocity by matching conditional velocities.
@@ 2.3 Optimal transport flows
A common and effective choice of the conditional probability paths is the optimal transport (OT) displacement map between the two Gaussians $p_{0}(\boldsymbol{x}|\boldsymbol{x}_{1})=\mathcal{N}(0,I)$ and $p_{1}(\boldsymbol{x}|\boldsymbol{x}_{1})=\mathcal{N}(\boldsymbol{x}_{1},\sigma_{min}I)$:
@@ 2.4 Variational parametrization
In Eijkelboom et al. [10], they show an alternative to directly matching the conditional velocity. They propose the following parametrization of the learned marginal velocity
$v_{t}^{\theta}(\boldsymbol{x}):=\int u_{t}(\boldsymbol{x}\mid\boldsymbol{x}_{1})q_{t}^{\theta}(\boldsymbol{x}_{1}\mid\boldsymbol{x})\mathrm{d}\boldsymbol{x}_{1}=\mathbb{E}_{\boldsymbol{x}_{1}\sim q_{t}^{\theta}(\boldsymbol{x}_{1}\mid\boldsymbol{x})}u_{t}(\boldsymbol{x}\mid\boldsymbol{x}_{1})$
Then the marginal velocity can be learned by matching a variational distribution $q_{t}^{\theta}(\boldsymbol{x}_{1}\mid\boldsymbol{x})$ to the ground truth $p_{t}(\boldsymbol{x}_{1}\mid\boldsymbol{x})$ by minimizing $D_{\text{KL}}(p_{t}(\boldsymbol{x}_{1}|\boldsymbol{x})\|q_{t}^{\theta}(\boldsymbol{x}_{1}|\boldsymbol{x}))$. In practice this is equivalent to maximum likelihood training:
$\mathcal{L}_{\mathrm{VFM}}(\theta)=-\mathbb{E}_{t\sim\mathcal{U}[0,1],\boldsymbol{x}_{1}\sim q(\boldsymbol{x}_{1}),\boldsymbol{x}\sim p_{t}(\boldsymbol{x}\mid\boldsymbol{x}_{1})}[\log q_{t}^{\theta}(\boldsymbol{x}_{1}\mid\boldsymbol{x})]$
@@
According to Eijkelboom et al. [10], this parametrization offers advantages when dealing with one-hot-encoded categorical variables. First, the generative paths become more realistic due to the inductive bias, which improves convergence by avoiding misaligned paths. Second, using cross-entropy loss instead of squared error enhances gradient behavior during training, thereby speeding up convergence.
@@ 3.1 Relational Data Generation
Relational databases are composed of multiple tables, where each table is a collection of records with a common structure. Tables may include one or more columns containing foreign keys, allowing records to refer to records in other tables. Consequently, a relational database can be represented as a graph, where nodes correspond to records and edges are defined by foreign-key relationships. In particular, the resulting graph is heterogeneous, since the nodes belong to different tables and have different types of features (i.e., the fields of a record). By convention, if a table $A$ contains foreign keys referring to records in table $B$, $A$ is called the child table and $B$ the parent table.
@@
Several approaches have been explored for relational data generation that differ in the order in which tables and topology are generated. An approach that has been recently proven effective [46] is to first generate the topology and then generate the tables one by one, conditioning on the topology and on previously generated tables:
@@
In this work we focus on the problem of conditional generation of the features $\mathbb{X}$ given the topology $\mathbb{G}$ of the relational database $p(\mathbb{X}\mid\mathbb{G})$. In this way the method for generating the foreign-key graph is independent from the method generating the features. This approach has advantages, as methods for generating large graphs are often quite various and different from deep generative models. Moreover, the complexity of our conditional generation method scales linearly with the size of the largest connected component, while most methods for graph generation scale quadratically [52].
@@ 3.2 Foreign-key Graph Generation
Since our work focuses on conditional generation of relational data content given a fixed topology, we adopt a simplified sampling approach for the foreign-key graph $\mathbb{G}$ (treating topology generation as orthogonal to our core contribution). For datasets where $\mathbb{G}$ has a large connected component, containing most or all nodes, we just keep the original topology $\mathbb{G}$. Otherwise, we build a new topology by sampling with replacement the connected components of $\mathbb{G}$ [20]. This method could be replaced by a dedicated graph sampler as in Xu et al. [46], which we consider an interesting direction for future work.
@@ 3.3 Generative Modeling from Single-Sample Data
We propose to learn a conditional generative model $p(\mathbb{X}\mid\mathbb{G})$ for the whole set of features $\mathbb{X}$, since in principle, it cannot always be decomposed into several independently and identically distributed (i.i.d.) samples, that in this case would correspond to the connected components of the graph. For example, in the movielens dataset [18], almost all records belong to the same connected component. Single-sample scenarios are common in large graph generation problems (e.g., social networks). Time series are another example where identifying i.i.d. samples is problematic. Nevertheless, successful modeling when disposing of only one sample remains feasible when the single sample is composed of weakly interacting components. This is the case of relational data, where records are usually not strongly dependent on all other records belonging to the same connected component. Moreover, the graph is sparse since the number of foreign keys is proportional to the number of records. Our method exploits structural regularities to enable effective learning from what is essentially a single sample.
@@
3.3.1 Graph-Conditional Flow Matching
In order to model $p(\mathbb{X}\mid\mathbb{G})$ using flow matching, we have to define the conditional flow $p_{t}(\mathbb{X}\mid\mathbb{X}_{1},\mathbb{G})$. We use the optimal transport conditional flow described in section 2 as conditional flow $p_{t}(\boldsymbol{x}^{i}\mid\boldsymbol{x}_{1}^{i})$, independent for each node $\boldsymbol{x}^{i}\in\mathbb{X}$ and for each component of a node (for each field of each record). We refer to the relative conditional velocity as $u_{t}(\mathbb{X}\mid\mathbb{X}_{1})$. This also holds for categorical components of features $\boldsymbol{x}^{i}$, which are encoded in continuous space using one-hot encodings [10]. Notice that this conditional flow does not depend on the topology $\mathbb{G}$, then we can refer to it as $p_{t}(\mathbb{X}\mid\mathbb{X}_{1})$. The objective is to learn the marginal velocity of the whole relational dataset $v_{t}(\mathbb{X}\mid\mathbb{G})$. We learn this using the variational parametrization discussed above:
@@ 3.3.1 Graph-Conditional Flow Matching
3.3.2 Variational Parametrization
We use a fully factorized distribution as variational approximation. This means that the distribution factorizes into an independent distribution for each component of each feature node $\boldsymbol{x}^{i}$. Every component $x^{k,d,j}$ of $\boldsymbol{x}^{i}$ is characterized by the table $k\in K$ it belongs to, containing $N_{k}$ records, and its column $d\in D_{k}$ where $D_{k}$ is the set of columns of table $k$. Then we can write the variational approximation as
$p_{\theta}(\mathbb{X}_{1}\mid\mathbb{X}_{t},\mathbb{G})=\prod_{k\in K}\prod_{d\in D_{k}}\prod_{j=1}^{N_{k}}p^{k,d}_{\theta}(x_{1}^{k,d,j}\mid\mathbb{X}_{t},\mathbb{G})$
@@ 3.3.2 Variational Parametrization
$p^{k,d}_{\theta}(x^{k,d,j}_{1}\mid\mathbb{X}_{t},\mathbb{G})=\mathrm{Categorical}\left(x^{k,d,j}_{1}\mid\mathbf{p}=f^{k,d}_{\theta_{2}}(x^{k,d,j}_{t},t,\varepsilon^{i}_{t})\right)$
@@
When component $x^{k,d,j}$ is a real number, we use the same architecture to parametrize the mean of a normal distribution with unitary variance:
$p_{\theta}(x^{k,d,j}_{1}\mid\mathbb{X}_{t},\mathbb{G})=\mathcal{N}(x^{k,d,j}_{1}\mid\mu=f^{k,d}_{\theta_{2}}(x^{k,d,j}_{t},t,\varepsilon^{i}_{t}),\sigma=1)$
In this case training corresponds to training a neural network regressor to predict $x^{k,d,j}_{1}$ using the squared error loss. We use a fixed unitary variance since learning the mean is sufficient to correctly parametrize the velocity field [10].
@@
$v^{\theta}_{t}(x^{k,d,j}\mid\mathbb{X}_{t},\mathbb{G})=\frac{\mathbb{E}_{x^{k,d,j}_{1}\sim p_{\theta}^{k,d}(x_{1}^{k,d,j}\mid\mathbb{X}_{t},\mathbb{G})}[x_{1}^{k,d,j}]-(1-\sigma_{\min})x^{k,d,j}}{1-(1-\sigma_{\min})t}$
@@
$v^{\theta}_{t}(x^{k,d,j}\mid\mathbb{X}_{t},\mathbb{G})=\frac{f^{k,d}_{\theta_{2}}(x^{k,d,j}_{t},t,\varepsilon^{i}_{t})-(1-\sigma_{\min})x^{k,d,j}}{1-(1-\sigma_{\min})t}$
@@
Notice that the neural network $f_{\theta_{2}}^{k,d}$ is specific to the column $d$ of table $k$. In particular, we use a different multi layer perceptron for each table $k$, with a prediction head (the last linear layer) for every component $d$. The inputs of $f_{\theta_{2}}^{k,d}$ are concatenated and flattened. Instead, the neural network $\eta_{\theta_{1}}$ computing node embeddings $\varepsilon^{i}_{t}=\eta_{\theta_{1}}(\mathbb{G},\mathbb{X}_{t})^{i}$, refers to a single GNN able to deal with heterogeneous graph data, i.e., graphs with multiple types of nodes, and as a consequence, different types of edges. We experimented with architectures based on GIN [47] and GATv2 [6]. In order to build a GNN compatible with heterogeneous graphs, we take an existing GNN model and use a dedicated GNN layer for each edge type, as showed in the documentation of the torch-geometric library [12]. When using GIN layers, we needed to embed node features in a common space. Figure 1 shows an overview of the denoiser’s architecture. More details are provided in Appendix B.
3.3.3 Implementation Details
@@
Following Kotelnikov et al. [27], during the preprocessing phase we transform continuous features using quantile transformation, so that all marginals of continuous features will be normally distributed. In order to handle missing data in numerical columns, we augment the tables including an auxiliary column containing binary variables indicating if the data is missing, and we fill missing values with the mean. For categorical columns with missing data, we simply include a new category "NaN". Tables that contain only foreign-key columns (and no features) are considered only when computing node embeddings. Time embedding is implemented according to Dhariwal and Nichol [9].
@@
In order to avoid overfitting, we randomly split nodes into a training and validation set, computing the loss only on training nodes. The experimental results report performance relative to models achieving the best validation loss during training. The training loss defined above implies full-batch training, or batches corresponding to connected components. The same applies during the generation phase. This means that it is necessary to fit entire connected components in memory. We observed relatively short training time: for the largest dataset, training took less than 20 minutes on a single GPU (see Appendix A for more details), for the other datasets just a few minutes.
@@
To solve the ODE generating the data, we used the Euler integration method with 100 steps. In our experiments the generation process was relatively fast: for the largest dataset we experimented with, the generation process took less than 10 seconds.

### v2

# Graph-Conditional Flow Matching for Relational Data Generation

@@ Abstract
Codehttps://github.com/DavideScassola/graph-conditional-flow-matching
@@ 1 Introduction
Tabular data synthesis (ctgan; shi2025tabdiff) has been subject to research for several years. Early methods were based on Bayesian networks (zhang2017privbayes), factor graphs (mckenna2019graphical) and autoregressive models (nowok2016synthpop). Most recent methods leverage the breakthroughs of deep-learning based generative models, from early latent variables models as VAEs (kingma2013auto) and GANs (goodfellow2014generative), to transformer-based autoregressive models (vaswani2017attention) and diffusion models (ho2020denoising). Despite the advancements in single table synthesis, generating multiple tables characterized by foreign-key constraints is a considerably more difficult task.
@@
Relational datasets can be represented as large graphs, where nodes correspond to records and edges denote foreign-key relationships. This introduces a dual challenge: (1) modeling potentially complex graph structures such as tables with multiple parents, or multiple types of relationships between two tables and (2) modeling statistical dependencies between records linked directly or indirectly through foreign keys.
@@
Relational data generation (patki2016synthetic; gueye2022row; solatorio2023realtabformergeneratingrealisticrelational; pang2024clavaddpm) is a less mature field, with few existing methods capable of properly handling complex database structures. In xu2022synthetic, they separately model the graph structure and then propose a method for generating the content of the relational database table-by-table. The generation of each record is conditioned on graph-derived node statistics and aggregated information from connected records. In concurrent work, hudovernik2024relational follow a similar approach, generating the content of the tables using a latent diffusion model conditioned on node embeddings produced by a separate model.
@@
In this work, we propose a method for generating the content of a relational dataset given the graph describing its structure. Inspired by recent advancements in image generation, we employ flow matching to train a flow-based generative model of the content of the entire relational dataset. In order to enable information propagation across connected records, the architecture of the learned denoiser includes a graph neural network (GNN) (scarselli2008graph; gnn2). This approach aims at maximizing expressiveness in modeling correlation between different records of the database, as information can be passed arbitrarily within a connected component through a GNN. Moreover, our framework is flexible as the conditioning graph can be complex, and scalable as we can generate large datasets.
@@
Using SyntheRela (hudovernik2024benchmarking), a recently developed benchmarking library, we prove the effectiveness of our method on several datasets, comparing it with several open-source approaches. Experimental results show our method achieves state-of-the-art performance in terms of fidelity of the generated data.
@@ 2.1 Continuous Normalizing Flows
Flow matching (lipman2022flow) is an emerging framework for training continuous normalizing flows (CNFs), a deep generative model that learns to transform random noise into target distributions through ordinary differential equations.
@@ 2.2 The Flow Matching Objective
Since direct access to the vector field $v_{t}$ of a CNF generating the data is unavailable, we cannot directly match a parameterized velocity field $v_{t}^{\theta}$ against the ground truth $v_{t}$. The main idea of flow matching is instead to define the underlying probability path as a mixture of conditional ”per-example” probability paths, that can be defined in a tractable way.
Let us denote by $p_{t}(\boldsymbol{x}\mid\boldsymbol{x}_{1})$ a conditional probability path such that $p_{0}(\boldsymbol{x}\mid\boldsymbol{x}_{1})=p_{0}(\boldsymbol{x})$ and $p_{1}(\boldsymbol{x}\mid\boldsymbol{x}_{1})$ is a distribution concentrated around $\boldsymbol{x}=\boldsymbol{x}_{1}$, as $p_{1}(\boldsymbol{x}\mid\boldsymbol{x}_{1})=\mathcal{N}(\boldsymbol{x}\mid\boldsymbol{x}_{1},\sigma^{2}I)$ with $\sigma$ small. We define the conditional velocity $u_{t}(\boldsymbol{x}\mid\boldsymbol{x}_{1})$ as the velocity generating the conditional probability path $p_{1}(\boldsymbol{x}\mid\boldsymbol{x}_{1})$. It is then possible to prove that the marginal velocity, defined as
$u_{t}(\boldsymbol{x})$ $=\int u_{t}(\boldsymbol{x}\mid\boldsymbol{x}_{1})p_{t}(\boldsymbol{x}_{1}\mid\boldsymbol{x})d\boldsymbol{x}_{1}$
$=\int u_{t}(\boldsymbol{x}\mid\boldsymbol{x}_{1})\frac{p_{t}(\boldsymbol{x}\mid\boldsymbol{x}_{1})q(\boldsymbol{x}_{1})}{p_{t}(\boldsymbol{x})}d\boldsymbol{x}_{1}$
@@
$\mathcal{L}_{\mathrm{FM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}[0,1],\boldsymbol{x}_{1}\sim q(\boldsymbol{x}_{1}),\boldsymbol{x}\sim p_{t}(\boldsymbol{x}\mid\boldsymbol{x}_{1})}\|v_{t}^{\theta}(\boldsymbol{x})-u_{t}(\boldsymbol{x}|\boldsymbol{x}_{1})\|^{2}$
known as the conditional flow matching objective, where a neural network learns to match the marginal velocity by matching conditional velocities.
@@ 2.3 Optimal Transport Flows
A common and effective choice of the conditional probability paths is the optimal transport (OT) displacement map between the two Gaussians $p_{0}(\boldsymbol{x}\mid\boldsymbol{x}_{1})=\mathcal{N}(0,I)$ and $p_{1}(\boldsymbol{x}\mid\boldsymbol{x}_{1})=\mathcal{N}(\boldsymbol{x}_{1},\sigma_{min}I)$:
@@ 2.4 Variational Parametrization
In eijkelboom2024variational, they show an alternative to directly matching the conditional velocity. They propose the following parametrization of the learned marginal velocity
$v_{t}^{\theta}(\boldsymbol{x})$ $=\int u_{t}(\boldsymbol{x}\mid\boldsymbol{x}_{1})q_{t}^{\theta}(\boldsymbol{x}_{1}\mid\boldsymbol{x})\mathrm{d}\boldsymbol{x}_{1}$
$=\mathbb{E}_{\boldsymbol{x}_{1}\sim q_{t}^{\theta}(\boldsymbol{x}_{1}\mid\boldsymbol{x})}u_{t}(\boldsymbol{x}\mid\boldsymbol{x}_{1})$
Then the marginal velocity can be learned by matching a variational distribution $q_{t}^{\theta}(\boldsymbol{x}_{1}\mid\boldsymbol{x})$ to the ground truth $p_{t}(\boldsymbol{x}_{1}\mid\boldsymbol{x})$ by minimizing $D_{\text{KL}}(p_{t}(\boldsymbol{x}_{1}\mid\boldsymbol{x})\|q_{t}^{\theta}(\boldsymbol{x}_{1}\mid\boldsymbol{x}))$. This is equivalent to maximum likelihood training:
$\mathcal{L}_{\mathrm{VFM}}(\theta)=-\mathbb{E}_{t\sim\mathcal{U}[0,1],\boldsymbol{x}_{1}\sim q(\boldsymbol{x}_{1}),\boldsymbol{x}\sim p_{t}(\boldsymbol{x}\mid\boldsymbol{x}_{1})}\log q_{t}^{\theta}(\boldsymbol{x}_{1}\mid\boldsymbol{x})$
@@
According to eijkelboom2024variational, this parametrization offers advantages when dealing with one-hot-encoded categorical variables. First, the generative paths become more realistic due to the inductive bias, which improves convergence by avoiding misaligned paths. Second, using cross-entropy loss instead of squared error enhances gradient behavior during training, thereby speeding up convergence.
Figure 1: Overview of the architecture of the denoiser for relational data. A relational dataset composed of multiple tables can be seen as a graph, where records are the nodes and foreign keys are the edges. The denoiser takes as input a relational dataset where noise was added to each record with noise level $t$. Firstly, a graph neural network (GNN) processes the entire graph and computes node embeddings $\varepsilon^{i}_{t}$ encoding context information for each record. Each record and its corresponding embedding are then processed independently by table-specific multi-layer perceptrons (MLPs), which predict the original clean records ($t=1$).
@@ 3.1 Relational Data Generation
Relational databases are composed of multiple tables, where each table is a collection of records with a common structure. Tables may include one or more columns containing foreign keys, allowing records to refer to records in other tables. Consequently, a relational database can be represented as a graph, where nodes correspond to records and edges are defined by foreign-key relationships. In particular, the resulting graph is heterogeneous, since the nodes belong to different tables and have different types of features (i.e., the fields of a record). By convention, if a table $A$ contains foreign keys referring to records in table $B$, AA is called the child table and $B$ the parent table.
@@
Several approaches have been explored for relational data generation that differ in the order in which tables and topology are generated. An approach that has been recently proven effective (xu2022synthetic) is to first generate the topology and then generate the tables one by one, conditioning on the topology and on previously generated tables:
@@
In this work we focus on the problem of conditional generation of the features $\mathbb{X}$ given the topology $\mathbb{G}$ of the relational database $p(\mathbb{X}\mid\mathbb{G})$. In this way, the method for generating the foreign-key graph is independent from the method generating the features. This modular approach has advantages, as methods for generating large graphs are often quite various and different from deep generative models. Moreover, the complexity of our conditional generation method scales linearly with the size of the graph, while most methods for graph generation scale quadratically (zhu2022survey). Finally, a conditional generation method is useful when one is interested in anonymizing the content but not the structure of a relational dataset, or when the topology is directly given by the user.
@@ 3.2 Foreign-key Graph Generation
Since our work focuses on conditional generation of relational data content given a fixed topology, we adopt a simplified sampling approach for the foreign-key graph $\mathbb{G}$ (treating topology generation as orthogonal to our core contribution). For datasets where $\mathbb{G}$ has a large connected component, containing most or all nodes, we just keep the original topology $\mathbb{G}$. Otherwise, we build a new topology by sampling with replacement the connected components of $\mathbb{G}$ (hudovernik2024relational). This method could be replaced by a dedicated graph sampler as in xu2022synthetic, which we consider an interesting direction for future work. We remark that the subject of this work is the conditional generation of the content given the foreign-key graph. An advantage of this approach is its modularity, as this method can be combined with any pure graph generation approach. Moreover, resampling the connected components has concrete applications, for instance, when all the observed topologies are well-represented in the training data.
@@ 3.3 Generative Modeling from Single-Sample Data
We propose to learn a conditional generative model $p(\mathbb{X}\mid\mathbb{G})$ for the whole set of features $\mathbb{X}$, since in principle, it cannot always be decomposed into several independently and identically distributed (i.i.d.) samples, that in this case would correspond to the connected components of the graph. For example, in the movielens dataset (movielens), almost all records belong to the same connected component. Single-sample scenarios are common in large graph generation problems (e.g., social networks). Time series are another example where identifying i.i.d. samples is problematic. Nevertheless, successful modeling when disposing of only one sample remains feasible when the single sample is composed of weakly interacting components. This is the case of relational data, where records are usually not strongly dependent on all other records belonging to the same connected component. Moreover, the graph is sparse since the number of foreign keys is proportional to the number of records. Our method exploits structural regularities to enable effective learning from what is essentially a single sample $\mathbb{D}=(\mathbb{X},\mathbb{G})$.
@@
Graph-Conditional Flow Matching
In order to model $p(\mathbb{X}\mid\mathbb{G})$ using flow matching, we have to define the conditional flow $p_{t}(\mathbb{X}\mid\mathbb{X}_{1},\mathbb{G})$. We use the optimal transport conditional flow described in Section 2 as conditional flow $p_{t}(\boldsymbol{x}^{i}\mid\boldsymbol{x}_{1}^{i})$, independent for each node $\boldsymbol{x}^{i}\in\mathbb{X}$ and for each component of a node (for each field of each record). We refer to the relative conditional velocity as $u_{t}(\mathbb{X}\mid\mathbb{X}_{1})$. This also holds for categorical components of features $\boldsymbol{x}^{i}$, which are encoded in continuous space using one-hot encodings (eijkelboom2024variational). Notice that this conditional flow does not depend on the topology $\mathbb{G}$, then we can refer to it as $p_{t}(\mathbb{X}\mid\mathbb{X}_{1})$. The objective is to learn the marginal velocity of the whole relational dataset $v_{t}(\mathbb{X}\mid\mathbb{G})$. We learn this using the variational parametrization discussed above:
@@
Variational Parametrization
Let $x^{k,d,j}$ be the $j_{th}$ value of column $d$ of table $k$, then we can write $\mathbb{X}=\left\{x^{k,d,j}\;\middle|\;k=1,\dots,K;\;d=1,\dots,D_{k};\;j=1,\dots,N_{k}\right\}$, where $D_{k}$ and $N_{k}$ are respectively the number of columns and records in table $k$. We use a fully factorized distribution as variational approximation. This means that the distribution factorizes into an independent distribution for each component $x^{k,d,j}$. Then we can write the variational approximation as
$p_{\theta}(\mathbb{X}_{1}\mid\mathbb{X}_{t},\mathbb{G})=\prod_{k=1}^{K}\prod_{d=1}^{D_{k}}\prod_{j=1}^{N_{k}}p^{k,d}_{\theta}(x_{1}^{k,d,j}\mid\mathbb{X}_{t},\mathbb{G})$
@@
$p^{k,d}_{\theta}(x^{k,d,j}_{1}\mid\mathbb{X}_{t},\mathbb{G})=\\ \mathrm{Categorical}\left(x^{k,d,j}_{1}\mid\mathbf{p}=f^{k,d}_{\theta_{2}}(x^{k,d,j}_{t},t,\varepsilon^{i}_{t})\right)$
@@
When component $x^{k,d,j}$ is a real number, we use the same architecture to parametrize the mean of a normal distribution with unit variance:
$p_{\theta}(x^{k,d,j}_{1}\mid\mathbb{X}_{t},\mathbb{G})=\\ \mathcal{N}(x^{k,d,j}_{1}\mid\mu=f^{k,d}_{\theta_{2}}(x^{k,d,j}_{t},t,\varepsilon^{i}_{t}),\sigma=1)$
In this case training corresponds to training a neural network regressor to predict $x^{k,d,j}_{1}$ using the squared error loss. We use a fixed unit variance since learning the mean is sufficient to correctly parametrize the velocity field (eijkelboom2024variational).
@@
$v^{\theta}_{t}(x^{k,d,j}\mid\mathbb{X}_{t},\mathbb{G})=\\ \frac{\mathbb{E}_{x^{k,d,j}_{1}\sim p_{\theta}^{k,d}(x_{1}^{k,d,j}\mid\mathbb{X}_{t},\mathbb{G})}[x_{1}^{k,d,j}]-(1-\sigma_{\min})x^{k,d,j}}{1-(1-\sigma_{\min})t}$
@@
$v^{\theta}_{t}(x^{k,d,j}\mid\mathbb{X}_{t},\mathbb{G})=\\ \frac{f^{k,d}_{\theta_{2}}(x^{k,d,j}_{t},t,\varepsilon^{i}_{t})-(1-\sigma_{\min})x^{k,d,j}}{1-(1-\sigma_{\min})t}$ (3)
Recall that $x^{k,d,j}$ can be either a continuous variable or a one-hot-encoded categorical variable. In this latter case, both $x^{k,d,j}$ and $v^{\theta}_{t}(x^{k,d,j}\mid\mathbb{X}_{t},\mathbb{G})$ are vectors.
@@
Notice that the neural network $f_{\theta_{2}}^{k,d}$ is specific to the column $d$ of table $k$. In particular, we use a different multi-layer perceptron for each table $k$, with a prediction head (the last linear layer) for every component $d$. The inputs of $f_{\theta_{2}}^{k,d}$ are concatenated and flattened.
Instead, the neural network $\eta_{\theta_{1}}$ computing node embeddings $\varepsilon^{i}_{t}=\eta_{\theta_{1}}(\mathbb{G},\mathbb{X}_{t})^{i}$, refers to a single GNN supporting heterogeneous graph data, i.e., graphs with multiple types of nodes, and as a consequence, different types of edges. We experimented with architectures based on GIN (xu2018powerful) and GATv2 (brody2021attentive). In order to build a GNN compatible with heterogeneous graphs, we take an existing GNN model and use a dedicated GNN layer for each edge type, as showed in the documentation of the torch-geometric library (pytorch_geometric_hetero). The messages produced by each edge-specific layer need to be of the same dimension, so that they can be summed together to obtain a node embedding. A detailed description of the employed GNN architectures is provided in Appendix B. Figure 1 shows an overview of the denoiser’s architecture.
Computational Complexity
Similarly to generative models for single-table data, the complexity of both training and generation of this method scales linearly with the number of records in the relational dataset. First of all, the computational complexity of each layer of the GNN scales linearly with the number of edges. In a relational dataset the total number of foreign keys is proportional to the number of records, since each record (node) has a fixed number of foreign keys (edges). Consequently, the GNN’s computational complexity is linear with the number of records. Second, the multi-layer perceptrons are applied independently to each record, meaning their computational complexity also scales linearly and can be efficiently parallelized.
Algorithm 1 Graph-Conditional Generation
Input: $\mathbb{G}\sim p(\mathbb{G})$
Parameter: Number of steps $T$
1: Sample $\mathbb{X}_{0}\sim\mathcal{N}(0,I)$
2:for $t=0$ to 11 step $\frac{1}{T}$ do
3:  $x_{t+\frac{1}{T}}^{k,d,j}=x_{t}^{k,d,j}+\frac{1}{T}v^{\theta}_{t}(x_{t}^{k,d,j}\mid\mathbb{X}_{t},\mathbb{G})$ Eq. (3)
4:end for
5:return $(\mathbb{X}_{1},\mathbb{G})$
Implementation Details
@@
Following kotelnikov2023tabddpm, during the preprocessing phase we transform continuous features using quantile transformation, so that all marginals of continuous features will be normally distributed. In order to handle missing data in numerical columns, we augment the tables including an auxiliary column containing binary variables indicating if the data is missing, and we fill missing values with the mean. This allows to preserve the information about missing data, and replicate missing data patterns in the generated samples. For categorical columns with missing data, we simply include a new category ”NaN”. Tables that contain only foreign-key columns (and no features) are considered only when computing node embeddings. Time embedding is implemented according to dhariwal2021diffusion.
@@
In order to avoid overfitting, we randomly split nodes into a training and validation set, computing the loss only on training nodes. The experimental results report performance relative to models achieving the best validation loss during training.
We performed full-batch training, as the entire dataset and the denoiser’s computations could be fitted into memory. This allowed us to denoise all record in parallel. For scenarios involving significantly larger datasets, where the graph or the GNN computations exceed available memory, one could employ mini-batching techniques for GNNs, or strategies for scaling to out-of-memory graphs.
We observed relatively short training time: for the largest dataset, training took less than 20 minutes on a single GPU (see Appendix A for more details), for the other datasets just a few minutes.
@@
To solve the ODE generating the data, we used the Euler integration method with 100 steps. In our experiments the generation process was relatively fast: for the largest dataset we experimented with, the generation process took less than 10 seconds. We summarize in Algorithm 1 the graph-conditional generation algorithm.